Distribution by ratio is common in partnership. First, describe how ratios function and then (give a numbers example) discuss how you would distribute profits and liquidated assets in ratio form.
Answer #1

Profit sharing ratio

In a partnership, Profit sharing ratio is agreed upon in the partnership agreement which is used to distribute profits & losses between the partners. In partnership agreement, Such ratio may be determined based on the contribution made by the each partner in the business or some other ratio may be specified. In case of no such partnership agreement is made, profits & losses are shared equally by the partners.

Example: Joe & Grace are partners in the business where the partnership agreement specifies the profit sharing ratio of 3:2.

At the end of the year, the profit of $20000 earned by the firm will be distributed in the ratio of 3:2.

Joe's share in profits (3/5) = $20000 x (3/5) = $12000

Grace's share in profits (2/5) = $20000 x (2/5) = $8000

Liquidation of firm

In the case of liquidation, following steps are followed:

1. Sale of non-cash asset

2. Division of gain or loss between partners in income ratio specified in the partnership agreement

3. Payment of external debts/liabilities

4. Distribution of remaining amount between partners in the capital ratio

Example 2

Continuing the data in Example 1, Joe & Grace have cash balance remaining in the business of $20000. Their capital balance in the books show Joe - $70000 Grace - $50000. They decided to liquidate the partnership, where they sold assets in the business having book value of $80000 for $85000 & made gain on such sale of $5000. External debt amounts to $35000.

Now, division of gain on sale of assets:-

Joe Grace
Capital balance 70000 50000
Division of gain of $5000 in 3:2 3000 3000
Capital balance after division of gain on sale of assets 73000 53000

Balance remaining after payment of external debt in the business

= Cash balance + Sale of Assets - External debt = $ (20000 + 85000 - 35000) = $70000

This balance will be distributed among Joe & Grace in their capital ratio after division of gain/loss on sale of assets.

Therefore, Joe's share = $70000 x ($73000 / $126000) = $40556

Grace's share = $70000 x ($53000 / $126000) = $29444.